express as the cofunction of a positive acute angle not using a calculator of tan -235 degrees.
Answer by venugopalramana(3286)

TAN(-235)=-TAN(235)SINCE IN IV QUADRANT TAN IS NEGATIVE....TAN(-X)=-TAN(X)
=-TAN(180+55)=-TAN55..SINCE IN III QUADRANT TAN IS POSITIVE..TAN(180+X)=TAN(X)